# Automatic Speech Recognition using Advanced Deep Learning Approaches: A survey

### v1

@@ 5.1 DTL and domain adaptation
MMD^2[f(X_S), f(X_T)] = ∥1nS∑_i=1^n_Sϕ(x_S_i) - 1nT∑_j=1^n_Tϕ(x_T_j)^2
where $\phi(\cdot)$ represents the mapping to a reproducing kernel Hilbert space (RKHS), and $n_{S}$, $n_{T}$ are the numbers of samples in the source and target domains, respectively.
@@ 5.2 FL-based
FL has significant potential for ASR systems, particularly in enhancing privacy and personalization. However, deploying this technology in ASR also comes with a set of challenges. Typically, in FL, data is inherently decentralized and can vary significantly across devices. This heterogeneity in speech data—due to differences in accents, dialects, languages, and background noise—can make it challenging to train a model that performs well across all nodes. Ensuring robustness and generalization of the ASR model under these conditions is a complex task. Moving on, FL requires periodic communication between the central server and the devices to update the model. For ASR systems, where models can be quite large, this can result in substantial communication overhead. Optimizing the efficiency of these updates, in terms of both bandwidth usage and energy consumption, especially on mobile devices, is a significant challenge. Besides, although FL is designed to enhance privacy by not sharing raw data, there are still privacy challenges. For instance, it’s possible to infer sensitive information from model updates. Ensuring that these updates do not leak private information about the users’ speech data is a critical concern that requires sophisticated privacy-preserving techniques like differential privacy or secure multi-party computation.
@@
Additionally, one of the advantages of FL is the ability to personalize models based on local data. However, balancing personalization with the need for a generally effective model—especially in a diverse ecosystem with varying speech patterns—is challenging. Achieving this balance without compromising the model’s overall performance or the personalization benefits is a key challenge. Moving forward, FL systems need to manage potentially thousands or millions of devices participating in the training process. Scalability issues, including managing updates from such a large and potentially unreliable network of devices, ensuring consistent model improvements, and handling devices joining or leaving the network, are significant technical hurdles. Lastly, in FL, the distribution of data across devices is often non-identically distributed (non-IID). This means that the speech data on one device might be very different from that on another, leading to challenges in training a model that generalizes well across all devices. Overcoming the bias introduced by non-IID data is a major challenge in FL for ASR.
5.3 RL-based
@@
Using RL in ASR systems offers promising avenues for improvement but also presents several challenges. Specifically, one of the primary challenges in applying RL to ASR is the issue of sparse and delayed rewards. In many ASR tasks, the system only receives feedback (rewards or penalties) after processing lengthy sequences of speech, making it difficult to attribute the reward to specific actions or decisions. This delay complicates the learning process, as the model struggles to identify which actions led to successful outcomes. Moreover, balancing exploration, trying new actions to discover their effects, with exploitation, using known actions that yield the best results, is a critical challenge in RL. In the context of ASR, this means the system must balance between adhering to known speech patterns and exploring new patterns or interpretations. Overemphasis on exploration can lead to inaccurate transcriptions, while excessive exploitation may prevent the model from adapting to new speakers or accents. Additionally, RL models typically require a significant amount of interaction data to learn effectively. In ASR, obtaining large volumes of labeled speech data, especially with user feedback, can be challenging and expensive. Additionally, RL algorithms can be sample-inefficient, meaning they need a lot of data before they start performing well, which can be a bottleneck in practical applications.
@@
Moving forward, most ASR systems are built using supervised learning techniques that rely on vast amounts of annotated data. Integrating RL into these systems poses technical challenges, as it requires a different training paradigm that focuses on learning from user interactions and feedback rather than static datasets. Besides, using RL in ASR often involves collecting and analyzing user feedback and interactions to improve the model. This raises concerns about user privacy and data security, as sensitive information might be inadvertently captured and used for training. Ensuring that data is handled securely and in compliance with privacy regulations is a significant challenge.
@@
Designing an appropriate reward function that accurately reflects the desired outcomes in ASR is challenging. The reward function must capture the nuances of speech recognition, such as accuracy, naturalness, and user satisfaction, which can be difficult to quantify. Poorly designed reward functions can lead to suboptimal learning outcomes or unintended behaviors. Lestly, ASR systems are used in a wide range of environments, from quiet offices to noisy streets. RL models need to adapt to these varying conditions, but training them to handle such diversity can be complex. The environment’s variability requires models that can generalize well across different acoustic conditions, which remains a challenge for RL-based ASR systems.
@@ 6.1 Personalized data augmentation for dysarthric and older people
While ASR technologies have advanced significantly, especially in recognizing typical speech patterns, they still struggle to accurately identify speech from individuals with dysarthria or older adults [120]. Gathering extensive datasets from these groups is challenging due to mobility limitations often associated with these populations. In this context, personalized data augmentation plays a crucial role [121, 122]. Personalized data augmentation tailores the training process to accommodate the unique speech patterns and challenges associated with these groups. Dysarthria, a motor speech disorder, and the natural aging process can lead to speech that deviates from the normative models typically used to train ASR systems, making accurate recognition difficult [123]. Personalized data augmentation introduces a wider range of speech variations into the training dataset, including those specific to dysarthric speakers or older adults. This can include variations in speech rate, pitch, articulation, and clarity. By training on this augmented dataset, the ASR system learns to recognize and accurately transcribe speech that exhibits these characteristics [124]. Moreover, this helps the ASR models generalize better to unseen examples of speech from dysarthric speakers or older adults. This enhanced generalization is crucial for real-world applications where the system encounters a wide range of speech variations. Moving forward, personalized data augmentation can employ specific techniques tailored to the needs of dysarthric speakers or older adults, such as simulating the slurring of words, varying speech tempo, or introducing background noise [125], commonly challenging for these groups [126]. Techniques like pitch perturbation, temporal stretching, and adding noise can simulate real-world conditions more accurately for these users. A typical example is presented in [127], where a unique approach utilizes speaker-dependent generative adversarial networks has been proposed.
@@ 6.2 Multitask learning for ASR
MTL enhances the performance of ASR systems by leveraging the inherent relatedness of multiple learning tasks to improve the generalization of the primary ASR task. This approach allows the ASR model to learn shared representations that capture underlying patterns across different but related tasks, leading to several key benefits [128]. Typically, MTL encourages the ASR system to learn representations that are beneficial across multiple tasks. This can lead to more robust feature extraction, as the model is not optimized solely for transcribing speech but also for other related tasks, such as speaker identification or emotion recognition. This shared learning process helps in capturing a broader range of speech characteristics, which can improve the ASR system’s ability to handle varied speech inputs. Moving one, By simultaneously learning related tasks, MTL acts as a form of regularization, reducing the risk of overfitting on the primary ASR task. This is because the model must find a solution that performs well across all tasks, preventing it from relying too heavily on noise or idiosyncrasies specific to the training data of the main task. Besides, learning auxiliary tasks alongside the main ASR task can improve the model’s generalization capabilities. For example, learning to identify the speaker or the language can provide additional contextual clues that help the ASR system better understand and transcribe ambiguous audio signals.
@@ 6.3 Federated multi-task learning and distillation for ASR
End-to-End ASR models Continuing to develop and refine end-to-end deep learning models that directly map speech inputs to text outputs without the need for intermediate representations (like phonetic transcriptions) [mamyrbayev2023hybrid]. Advances in models such as Transformer and Conformer architectures, which can capture long-range dependencies in speech, are promising for improving ASR accuracy and efficiency [liu2023sfa]. Specifically, end-to-end models in ASR aim to directly convert speech input into text output using a single neural network architecture. This simplifies the ASR pipeline by bypassing traditional stages such as acoustic, pronunciation, and language modeling [yang2023attention].
An end-to-end ASR model maps an input sequence of speech features $X=(x_{1},x_{2},\ldots,x_{T})$ to an output sequence of tokens $Y=(y_{1},y_{2},\ldots,y_{N})$. The model function $f$ with parameters $\theta$ aims to minimize the difference between the predicted sequence and the ground truth. Two common loss functions are used [almadhor2023e2e]:
  • CTC Loss: Connectionist Temporal Classification (CTC) introduces a ’blank’ token for aligning sequences. The CTC loss is given by:
    $\mathcal{L}_{\text{CTC}}(\theta)=-\sum_{(X,Y)\in\mathcal{D}}\log P(Y|X;\theta)$
    where $\mathcal{D}$ is the dataset, and $P(Y|X;\theta)$ is the sequence probability.
  • Attention-based Seq2Seq Loss: This uses an encoder-decoder architecture with attention to predict each token. The loss is the negative log-likelihood:
    $\mathcal{L}_{\text{Seq2Seq}}(\theta)=-\sum_{(X,Y)\in\mathcal{D}}\log P(Y|X;\theta)$
Besides, model architectures can be described as follows:
  • Recurrent Neural Networks (RNNs) including Long Short-Term Memory (LSTM) networks for temporal dependencies [vander2023using].
  • Convolutional Neural Networks (CNNs) for capturing local patterns in speech features.
  • Transformer Models utilize self-attention mechanisms for parallel processing and superior performance [wang2022optimizing].
Gradient-based methods like stochastic gradient descent (SGD) or Adam are used to minimize the loss function, optimizing the model parameters $\theta$ to improve the speech-to-text mapping. However, end-to-end models face challenges such as data scarcity for low-resource languages, computational resource demands, and integrating external language models for linguistic improvements [qu2023emphasizing].
6.4 Domain-specific language modeling
[…]
$P(w_{n}|w_{n-1},w_{n-2},\ldots,w_{n-(n-1)})=\frac{C(w_{n-(n-1)},\ldots,w_{n})}{C(w_{n-(n-1)},\ldots,w_{n-1})}$ (8)
Additionally, data augmentation plays a crucial role in ASR-based DA. Typically, data augmentation in domain-specific ASR involves generating synthetic training examples by altering existing recordings or using text-to-speech (TTS) systems to create new audio samples from domain-specific texts. Mathematically, augmentation techniques can include time stretching, pitch shifting, adding noise, or simulating room acoustics, each represented as transformations $T$ applied to the original audio signal $x(t)$:
$x^{\prime}(t)=T(x(t);\phi)$ (9)
where $\phi$ represents the parameters of the transformation. Moving on, regularization techniques are applied to prevent overfitting to the domain-specific data, ensuring the model maintains generalization capabilities. Techniques such as dropout, L1/L2 regularization, or elastic net involve adding terms to the loss function or modifying the optimization process to penalize large weights or complex models. For instance, L2 regularization adds a penalty equal to the square of the magnitude of coefficients:
$L^{\prime}(D;\theta)=L(D;\theta)+\lambda\|\theta\|^{2}$ (10)
where $\lambda$ is a regularization parameter.
While DSLM offers substantial benefits, challenges remain, such as data scarcity in niche domains, the need for continual model updates to keep pace with evolving language use, and ensuring privacy and security in sensitive domains like healthcare and finance. Addressing these challenges through innovative approaches in model training, data augmentation, and privacy-preserving technologies will be crucial for the advancement of domain-specific ASR systems.
Privacy preservation With the advancements in ML, DL, and DTL, ASR systems have become more practical and scalable. However, these systems also introduce serious privacy concerns due to the abundance of sensitive acoustic and textual information in speech data.
While open-source and offline ASR systems can mitigate privacy risks, online DTL-based systems can amplify these threats. Additionally, the transcription performance of offline and open-source ASR systems is typically inferior to cloud-based ASR systems, especially in real-world scenarios [ahmed2020preech]. In this context, the $D_{S}$ data may contain sensitive information that needs to be protected. Thus, preserving users’ privacy during the knowledge transfer from the source domain (SD) to the target domain (TD) becomes a critical issue.
To address this challenge, future research efforts should focus on integrating effective security and privacy protection strategies. Examples include decentralized DTL approaches utilizing blockchain technology [ul2020decentralized, wang2021enabling] and federated DTL methods [zhang2021federated, maurya2021federated]. These approaches aim to ensure privacy-preserving knowledge transfer in DTL-based ASR systems.
6.5 Interpretation of DTL models
DTL-based ASR models, despite their success, are often considered as "black box" systems lacking interpretability. This lack of interpretability raises doubts about the credibility and repeatability of their decisions, making it crucial to explain the reasoning behind their predictions. The field of explainable and interpretable ML/DL has gained increasing interest in various applications, including speech processing.
Several studies have explored the interpretability of DTL models. In one study [ramakrishnan2016towards], an agent was designed to explain how it learns a new task using prior common knowledge, aiming to enhance users’ trust and acceptance of the system results. Another study [kim2019the] defined interpretable features in a DTL algorithm and examined the relationship between the SD and TD in the task, focusing on the interpretability of the pretrained DTL model.
The work by Lee et al. [lee2021interpretable] introduced a knowledge distillation approach that generated interpretable embedding procedure (IEP) knowledge based on PCA and transferred it to the student network using a message passing neural network. This approach enhanced interpretability while maintaining accuracy through multi-task learning. Additionally, Carr et al. proposed an interpretable staged TL (iSTL) scheme for accurate and explainable classification of optical coherence tomography (OCT) scans with a small sample size [carr2021interpretable]. iSTL outperformed DTL techniques on unseen data, utilizing clinical features for predictions with interpretable attention maps.
@@ 6.5 Interpretation of DTL models
Delving deeper into techniques for FL distillation (optimizing model size) within FL frameworks is essential. This exploration involves researching methods to compress neural ASR models effectively while ensuring their performance remains intact, especially tailored for edge devices with storage and computational constraints. It is imperative to investigate the trade-offs associated with reducing model size while maintaining performance metrics like WER. Developing strategies to strike a balance between downsizing models and preserving satisfactory performance levels within federated learning environments is crucial.
6.6 Recent RL techniques for ASR
@@
Exploring incremental DRL approaches [136, 137, 138] in DRL-based ASR systems, could be very interesting. This approach involves the model continuously learning from newly acquired data and dynamically adjusting its ASR functionalities over time. By incrementally updating its knowledge base, the model can enhance its performance without necessitating full retraining, thus enabling continual enhancement of ASR systems. This capability not only fosters greater resilience and adaptability in speech recognition capabilities but also offers potential applications in scenarios where real-time adaptation to changing conditions is crucial, such as in noisy environments or with varying speaker accents. Moreover, incremental RL can potentially lead to more efficient use of computational resources, as the model only needs to focus on learning from new data, rather than reprocessing the entire dataset. Further research in this area could unlock new possibilities for ASR systems to evolve and improve over time, ultimately enhancing their usability and effectiveness in diverse real-world settings.
@@
Although some ASR schemes based on RL have been proposed, there remains a notable scarcity in the application of RL techniques to enhance ASR methods. While policy gradient and Q-learning are commonly employed, the realm of RL encompasses various subcategories such as double deep Q-network (DDQN), actor-critic (AC), deep deterministic policy gradien (DDPG), and more [139], which hold promise for advancing ASR with innovative approaches. Researchers are encouraged to delve into these diverse DRL-based methods to further enrich the field of ASR for both AM and LM fields.
@@ 6.7 Online DTL
Online DTL combines the principles of TL and online learning with deep neural networks, enabling models to adapt in real-time to new tasks or data distributions. This approach is beneficial in dynamic environments where data arrives sequentially. DL models, specifically Deep Neural Networks (DNNs), learn through optimizing the weights $\theta$ to minimize a loss function $L$, which measures the discrepancy between predicted outputs $\hat{y}$ and true outputs $y$: $\theta^{*}=\arg\min_{\theta}L(D;\theta)$. On the other hand, TL improves learning in a new target task through the transfer of knowledge from a related source task, adapting a pre-trained model $\theta_{S}$ on $D_{S}$ to a $D_{T}$: $\theta_{T}^{*}=\arg\min_{\theta}L(D_{T};\theta_{T})$. In this regard, online learning updates the model incrementally as new data $(x_{t},y_{t})$ arrives:
@@
Where $\alpha_{t}$ is the learning rate, and $\nabla_{\theta}L$ is the gradient of the loss with respect to $\theta$. Moving on, online TL integrates these concepts to continuously adapt a deep learning model to new tasks or data streams, often involving techniques such as feature extraction, fine-tuning, model adaptation, and continual learning. The adaptation process at each time step $t$ can be viewed as $\theta_{t+1}^{*}=\arg\min_{\theta}L(D_{T_{t}};\theta_{T_{t}})$. Where $D_{T_{t}}$ represents the data available at time $t$, including new target domain data.
The approach of online DTL offers a forward-looking solution to this issue. Particularly, discrepancies in class distributions and the representation of features between the $D_{S}$ and $D_{T}$ amplify the complexity of online DTL [140]. To navigate the complexities mentioned, online DTL has been dissected into two primary methodologies. The first, known as homogeneous online DTL, operates on the premise of a unified feature space across both domains. Conversely, heterogeneous online DTL acknowledges the distinct feature spaces intrinsic to each domain [141]. An exemplary solution to the challenges of heterogeneous online DTL includes leveraging unlabeled instances of co-occurrence to forge a connective bridge between the $D_{S}$ and $D_{T}$, facilitating the precursor to knowledge transfer [142]. Furthering the discourse, online DTL augmented with extreme learning machines introduces a novel framework [143]. Addressing the challenge of limited data in the $D_{T}$, the technique of transfer learning with lag (TLL), rooted in shallow neural network embeddings, has been applied. This method ensures the continuity of knowledge transfer, notwithstanding fluctuations in the feature set.

### v2

@@ 4.3 FL-based
FL has significant potential for DTL systems, particularly in enhancing privacy and personalization. However, deploying this technology in DTL also comes with a set of challenges. Typically, in FL, data is inherently decentralized and can vary significantly across devices. This heterogeneity in speech data—due to differences in accents, dialects, languages, and background noise—can make it challenging to train a model that performs well across all nodes. Ensuring robustness and generalization of the DTL model under these conditions is a complex task. Moving on, FL requires periodic communication between the central server and the devices to update the model. For DTL systems, where models can be quite large, this can result in substantial communication overhead. Optimizing the efficiency of these updates, in terms of both bandwidth usage and energy consumption, especially on mobile devices, is a significant challenge. Besides, although FL is designed to enhance privacy by not sharing raw data, there are still privacy challenges. For instance, it’s possible to infer sensitive information from model updates. Ensuring that these updates do not leak private information about the users’ speech data is a critical concern that requires sophisticated privacy-preserving techniques like differential privacy or secure multi-party computation.
@@
Additionally, one of the advantages of FL is the ability to personalize models based on local data. However, balancing personalization with the need for a generally effective model—especially in a diverse ecosystem with varying speech patterns—is challenging. Achieving this balance without compromising the model’s overall performance or the personalization benefits is a key challenge. Moving forward, FL systems need to manage potentially thousands or millions of devices participating in the training process. Scalability issues, including managing updates from such a large and potentially unreliable network of devices, ensuring consistent model improvements, and handling devices joining or leaving the network, are significant technical hurdles. Lastly, in FL, the distribution of data across devices is often non-IID. This means that the speech data on one device might be very different from that on another, leading to challenges in training a model that generalizes well across all devices. Overcoming the bias introduced by non-IID data is a major challenge in FL for ASR.
4.4 DRL-based
@@
Using DRL in ASR systems offers promising avenues for improvement but also presents several challenges. Specifically, one of the primary challenges in applying DRL to ASR is the issue of sparse and delayed rewards. In many ASR tasks, the system only receives feedback (rewards or penalties) after processing lengthy sequences of speech, making it difficult to attribute the reward to specific actions or decisions. This delay complicates the learning process, as the model struggles to identify which actions led to successful outcomes. Moreover, balancing exploration, trying new actions to discover their effects, with exploitation, using known actions that yield the best results, is a critical challenge in DRL. In the context of ASR, this means the system must balance between adhering to known speech patterns and exploring new patterns or interpretations. Overemphasis on exploration can lead to inaccurate transcriptions, while excessive exploitation may prevent the model from adapting to new speakers or accents. Additionally, DRL models typically require a significant amount of interaction data to learn effectively. In ASR, obtaining large volumes of labeled speech data, especially with user feedback, can be challenging and expensive. Additionally, DRL algorithms can be sample-inefficient, meaning they need a lot of data before they start performing well, which can be a bottleneck in practical applications.
@@
Moving forward, most ASR systems are built using supervised learning techniques that rely on vast amounts of annotated data. Integrating DRL into these systems poses technical challenges, as it requires a different training paradigm that focuses on learning from user interactions and feedback rather than static datasets. Besides, using DRL in ASR often involves collecting and analyzing user feedback and interactions to improve the model. This raises concerns about user privacy and data security, as sensitive information might be inadvertently captured and used for training. Ensuring that data is handled securely and in compliance with privacy regulations is a significant challenge.
@@
Designing an appropriate reward function that accurately reflects the desired outcomes in ASR is challenging. The reward function must capture the nuances of speech recognition, such as accuracy, naturalness, and user satisfaction, which can be difficult to quantify. Poorly designed reward functions can lead to suboptimal learning outcomes or unintended behaviors. Lastly, ASR systems are used in a wide range of environments, from quiet offices to noisy streets. DRL models need to adapt to these varying conditions, but training them to handle such diversity can be complex. The environment’s variability requires models that can generalize well across different acoustic conditions, which remains a challenge for DRL-based ASR systems.
@@ 5.1 Personalized data augmentation for dysarthric and older people
While DTL technologies have advanced significantly, especially in recognizing typical speech patterns, they still struggle to accurately identify speech from individuals with dysarthria or older adults [121]. Gathering extensive datasets from these groups is challenging due to mobility limitations often associated with these populations. In this context, personalized data augmentation plays a crucial role [122, 123]. Personalized data augmentation tailors the training process to accommodate the unique speech patterns and challenges associated with these groups. Dysarthria, a motor speech disorder, and the natural aging process can lead to speech that deviates from the normative models typically used to train DTL systems, making accurate recognition difficult [124]. Personalized data augmentation introduces a wider range of speech variations into the training dataset, including those specific to dysarthric speakers or older adults. This can include variations in speech rate, pitch, articulation, and clarity. By training on this augmented dataset, the DTL system learns to recognize and accurately transcribe speech that exhibits these characteristics [125]. Moreover, this helps the DTL models generalize better to unseen examples of speech from dysarthric speakers or older adults. This enhanced generalization is crucial for real-world applications where the system encounters a wide range of speech variations. Moving forward, personalized data augmentation can employ specific techniques tailored to the needs of dysarthric speakers or older adults, such as simulating the slurring of words, varying speech tempo, or introducing background noise [126], commonly challenging for these groups [127]. Techniques like pitch perturbation, temporal stretching, and adding noise can simulate real-world conditions more accurately for these users. A typical example is presented in [128], where a unique approach utilizes speaker-dependent generative adversarial networks has been proposed.
@@ 5.2 Multitask learning for ASR
MTL enhances the performance of DTL systems by leveraging the inherent relatedness of multiple learning tasks to improve the generalization of the primary ASR task. This approach allows the ASR model to learn shared representations that capture underlying patterns across different but related tasks, leading to several key benefits [129]. Typically, multitask learning (MTL) encourages the ASR system to learn representations that are beneficial across multiple tasks. This can lead to more robust feature extraction, as the model is not optimized solely for transcribing speech but also for other related tasks, such as speaker identification or emotion recognition. This shared learning process helps in capturing a broader range of speech characteristics, which can improve the ASR system’s ability to handle varied speech inputs. Moving on, by simultaneously learning related tasks, MTL acts as a form of regularization, reducing the risk of overfitting on the primary ASR task. This is because the model must find a solution that performs well across all tasks, preventing it from relying too heavily on noise or idiosyncrasies specific to the training data of the main task. Besides, learning auxiliary tasks alongside the main ASR task can improve the model’s generalization capabilities. For example, learning to identify the speaker or the language can provide additional contextual clues that help the ASR system better understand and transcribe ambiguous audio signals.
@@ 5.3 Federated multi-task learning and distillation for ASR
Delving deeper into techniques for FL distillation (optimizing model size) within FL frameworks is essential. This exploration involves researching methods to compress neural ASR models effectively while ensuring their performance remains intact, especially tailored for edge devices with storage and computational constraints. It is imperative to investigate the trade-offs associated with reducing model size while maintaining performance metrics like WER. Developing strategies to strike a balance between downsizing models and preserving satisfactory performance levels within FL environments is crucial.
5.4 Recent DRL techniques for ASR
@@
Exploring incremental DRL approaches [137, 138, 139] in DRL-based ASR systems, could be very interesting. This approach involves the model continuously learning from newly acquired data and dynamically adjusting its ASR functionalities over time. By incrementally updating its knowledge base, the model can enhance its performance without necessitating full retraining, thus enabling continual enhancement of ASR systems. This capability not only fosters greater resilience and adaptability in speech recognition capabilities but also offers potential applications in scenarios where real-time adaptation to changing conditions is crucial, such as in noisy environments or with varying speaker accents. Moreover, incremental DRL can potentially lead to more efficient use of computational resources, as the model only needs to focus on learning from new data, rather than reprocessing the entire dataset. Further research in this area could unlock new possibilities for ASR systems to evolve and improve over time, ultimately enhancing their usability and effectiveness in diverse real-world settings.
@@
Although some DTL schemes based on DRL have been proposed, there remains a notable scarcity in the application of DRL techniques to enhance DTL methods. While policy gradient and Q-learning are commonly employed, the realm of DRL encompasses various subcategories such as double deep Q-network (DDQN), actor-critic (AC), deep deterministic policy gradien (DDPG), and more [140], which hold promise for advancing DTL with innovative approaches. Researchers are encouraged to delve into these diverse DRL-based methods to further enrich the field of DTL for both AM and LM fields.
@@ 5.5 Online DTL
Online DTL combines the principles of DTL and online learning with DNNs, enabling models to adapt in real-time to new tasks or data distributions. This approach is beneficial in dynamic environments where data arrives sequentially. DL models, specifically DNNs, learn through optimizing the weights $\theta$ to minimize a loss function $L$, which measures the discrepancy between predicted outputs $\hat{y}$ and true outputs $y$: $\theta^{*}=\arg\min_{\theta}L(D;\theta)$. On the other hand, DTL improves learning in a new target task through the transfer of knowledge from a related source task, adapting a pre-trained model $\theta_{S}$ on $D_{S}$ to a $D_{T}$: $\theta_{T}^{*}=\arg\min_{\theta}L(D_{T};\theta_{T})$. In this regard, online learning updates the model incrementally as new data $(x_{t},y_{t})$ arrives:
@@
Where $\alpha_{t}$ is the learning rate, and $\nabla_{\theta}L$ is the gradient of the loss with respect to $\theta$. Moving on, online DTL integrates these concepts to continuously adapt a deep learning model to new tasks or data streams, often involving techniques such as feature extraction, fine-tuning, model adaptation, and continual learning. The adaptation process at each time step $t$ can be viewed as $\theta_{t+1}^{*}=\arg\min_{\theta}L(D_{T_{t}};\theta_{T_{t}})$. Where $D_{T_{t}}$ represents the data available at time $t$, including new target domain data.
The approach of online DTL offers a forward-looking solution to this issue. Particularly, discrepancies in class distributions and the representation of features between the $D_{S}$ and $D_{T}$ amplify the complexity of online DTL [141]. To navigate the complexities mentioned, online DTL has been dissected into two primary methodologies. The first, known as homogeneous online DTL, operates on the premise of a unified feature space across both domains. Conversely, heterogeneous online DTL acknowledges the distinct feature spaces intrinsic to each domain [142]. An exemplary solution to the challenges of heterogeneous online DTL includes leveraging unlabeled instances of co-occurrence to forge a connective bridge between the $D_{S}$ and $D_{T}$, facilitating the precursor to knowledge transfer [143]. Furthering the discourse, online DTL augmented with extreme learning machines introduces a novel framework [144]. Addressing the challenge of limited data in the $D_{T}$, the technique of DTL with lag, rooted in shallow neural network embeddings, has been applied. This method ensures the continuity of knowledge transfer, notwithstanding fluctuations in the feature set.